### v1

# Impact of Sunglasses on One-to-Many Facial Identification Accuracy

Abstract
One-to-many facial identification is documented to achieve high accuracy in the case where both the probe and the gallery are ‘mugshot quality’ images. However, an increasing number of documented instances of wrongful arrest following one-to-many facial identification have raised questions about its accuracy. Probe images used in one-to-many facial identification are often cropped from frames of surveillance video and deviate from ‘mugshot quality’ in various ways. This paper systematically explores how the accuracy of one-to-many facial identification is degraded by the person in the probe image choosing to wear dark sunglasses. We show that sunglasses degrade accuracy for mugshot-quality images by an amount similar to strong blur or noticeably lower resolution. Further, we demonstrate that the combination of sunglasses with blur or lower resolution results in even more pronounced loss in accuracy. These results have important implications for developing objective criteria to qualify a probe image for the level of accuracy to be expected if it used for one-to-many identification. To ameliorate the accuracy degradation caused by dark sunglasses, we show that it is possible to recover about 38% of the lost accuracy by synthetically adding sunglasses to all the gallery images, without model re-training. We also show that increasing the representation of wearing-sunglasses images in the training set can largely reduce the error rate. The image set assembled for this research will be made available to support replication and further research into this problem.
I Introduction
One-to-many facial identification has attracted significant attention due to it being used in a growing number of cases of wrongful arrest [8, 16, 29, 37]. In real-world scenarios, probe images used in one-to-many facial identification are often extracted from surveillance video. Such images may suffer from quality issues such as blur, low resolution, off-angle pose, and the presence of occlusions such as sunglasses and / or hats. Recent work has systematically investigated how image-based factors like resolution and blur degrade the accuracy of one-to-many identification [24]. However, subject-based factors such as wearing sunglasses or caps, or extreme pitch angles, which may be common in surveillance camera images and may substantially degrade identification accuracy, have not been thoroughly addressed.
Figure 1: How much does wearing sunglasses degrade one-to-many facial identification accuracy? One-to-many identification has high accuracy when the probe and gallery images are all high quality [10]. This work explores how accuracy degrades when the person in the probe image wears sunglasses, and alternatives for ameliorating this problem.
This paper presents the systematic exploration of the effect of sunglasses on the accuracy of one-to-many facial identification. First, to create a research dataset for this work, we assemble a test set, named “ND-sunglasses”, that contains 16,350 high-quality image pairs (original and with sunglasses), where the only difference in each pair of images is the sunglasses region. Then, we sample a range of degrees of blur and lower resolutions to select ones that have a similar-sized effect on accuracy as sunglasses. Then, we systematically analyze the impact of the combination of sunglasses, blur, and low resolution on accuracy. The results reveal a pronounced accuracy degradation, showing the importance of this study. Next, we introduce a simple but effective approach that can recover up to 37.63% of the accuracy lost, without model re-training, by adding sunglasses to the gallery images. Lastly, we also show that increasing the representation of wearing-sunglasses images in the training set can largely reduce the error rate.
Our study contributes to a better understanding of the accuracy that can be expected from one-to-many facial identification with probe images taken from surveillance video. The accuracy is far lower than has been documented for mugshot-quality images [10, 9, 19], and a variety of different factors can contribute to degrading the accuracy.
Contributions of this work include the following:
  • Experimental results investigated the impact of adding sunglasses to images in a one-to-many identification scenario. Our findings indicate that probe images with sunglasses exhibited significantly reduced accuracy in one-to-many identification.
  • Analyzing the effect of stacking blur and low resolution on the impact of sunglasses in probe images. In real-world surveillance camera images, a combination of these image quality factors is commonly encountered.
  • Results showing that when a probe image contained sunglasses while the gallery images did not, adding sunglasses to the gallery images can mitigate the accuracy loss due to sunglasses in the probe image.
  • Increasing the representation of sunglasses-wearing images in the training dataset substantially reduces error rates. This finding offers a practical approach to mitigating the accuracy loss associated with sunglasses.
  • Releasing the first dataset focused on studying the effect of sunglasses in one-to-many facial identification to enable replication and extension of our results.
II Related Work
Figure 2: The pipeline for creating sunglasses-added versions of original probe images. An Android emulator and ManyCam are used to handle the image input and processing steps, the FaceLab app adds sunglasses to the images, and screenshot the FaceLab results to store the images.
The recent NIST report [10] stands as the largest experimental investigation of the accuracy of one-to-many facial identification, given the large number of matchers evaluated and the size of image datasets used. The report emphasizes that demographic differences in the false positive identification rate (FPIR) are significantly larger than those in the false negative identification rate (FNIR), with FPIR differences potentially reaching one or two orders of magnitude (i.e., 10x, 100x). In contrast, FNIR differences are smaller and more specific to the algorithms used. Krishnapriya et al.[19] also examined one-to-many matching accuracy, utilizing the MORPH [26] dataset. Their findings agree with the NIST study, that when the person in the probe has an enrolled image, the one-to-many search is highly likely to return the correct identity. Both [10] and [19] are clear to state that, if the person in the probe image does not have an image enrolled in the gallery, any result returned by one-to-many facial identification must be a false positive identification.
Another NIST report [9] discusses the impressive accuracy achieved when matching mugshot-quality probe images to mugshot-quality enrolled images – “… in galleries containing 12 million individuals, with rank one miss rates approaching 0.1%.” (The term “mugshot quality” refers generally to face images with frontal pose, neutral expression and controlled lighting and background, as also seen with driver’s licence or visa / passport images.) The NIST report notes that the primary datasets comprise good quality images, which is an important factor in the low false positive identification rates achieved by the algorithms. However, this NIST report also notes that, “Recognition in other circumstances, particularly those without a dedicated photographic environment and human or automated quality control checks, will lead to declines in accuracy.”
One factor in the documented instances of wrongful arrest following one-to-many facial identification [8, 16, 29, 37] appears to be a mistaken confidence in the accuracy that can be achieved when the probe image is surveillance camera quality rather than mugshot quality. This has led to work to document how factors that can occur in face images from surveillance video degrade the accuracy of one-to-many facial identification.
Bhatta et al. [24] investigate how image quality factors such as blur and reduced resolution affect of one-to-many search using the MORPH dataset. They report that increased blur in the probe image, or reduced resolution of the face in the probe image, can significantly increase the false positive identification rate. They also find a larger accuracy degradation for females than for males.
In addition to image-based factors such as blur and resolution, face images taken from surveillance video may capture a person wearing dark sunglasses. Existing face attribute editing methods could potentially be used to add sunglasses to a face image [13, 20, 11, 28]. However, these methods suffer from unsatisfactory performance on high-resolution images. Moreover, due to the lack of sunglasses labels in popular existing training sets for GANs, adding sunglasses to an image is not (yet) a stable operation with these approaches. Hence, we choose to use a commercial (non-GAN) tool to reliably add realistic-looking sunglasses without changing other regions of the image, and so in an identity-preserving way.
[7] is the most related work, which proposes an occlusion dataset for one-to-many research. Whereas, this dataset has only 161 probe identities and no gender and race labels to address the possible ethical bias, which preventing us from obtaining statistically meaningful results.
This paper extends recent work in exploring the degradation in one-to-many facial identification accuracy caused by wearing sunglasses and providing two possible solutions to mitigate this degradation.
III Dataset
Figure 3: Examples of the original and synthetic sunglasses-added images. Adding synthetic sunglasses alters only the sunglasses region of the original image. Note that the style of the synthetic sunglasses vary between images.
There are specific requirements for a dataset suitable for exploring the effect of sunglasses in probe images on the accuracy of one-to-many identification. One, similar quality images of the same persons, wearing sunglasses in one image and not in another, are required for use as probe images for comparison experiments. Two, additional images of the same persons are required for use as the enrolled gallery images. The additional gallery images of a person should be acquired in a different session than the probe images, to avoid any “same-session effect” in accuracy estimates.
There are existing test sets of face images covering a variety of aspects of facial recognition, including pose [38, 27], age difference [39, 23], size of face in image [5], in-the-wild images and video [15, 18, 31, 22], and illumination and facial hair [35]. None of these focus specifically on the effects of sunglasses. The IJB datasets [18, 31, 22] do contain images of some persons wearing sunglasses. However, the IJB datasets emphasize in-the-wild imagery, which would be appropriate for probe image in our experiments, but not enrolled gallery images. Also, the presence of similar quality images of the same person, both wearing and not wearing sunglasses, would happen only by accident in the IJB datasets. Additionally, NIST has discontinued distribution of the IJB datasets. [^1] https://www.nist.gov/itl/iad/ig/ijb-c-dataset-request-form The AR face image dataset [21] introduced by Martinez in 1998 contains images of persons wearing sunglasses. However, there are only 126 persons in the dataset, the images a grayscale rather than color, and all persons wear the same model/style of sunglasses, images were acquired in two sessions per person, meaning that each person could have only one gallery image in one-to-many identification experiments. In the end, we could not identify an existing dataset suitable for exploring the effects of wearing sunglasses in the context of one-to-many facial identification.
Figure 4: Examples of FaceLab’s synthetic sunglasses on AR dataset. This 2x4 matrix displays two subjects (rows) under four conditions (columns): normal expression, smiling expression, real sunglasses, and synthetic sunglasses applied to the normal expression.
To assemble a suitable dataset for one-to-many identification experiments, we started with the Notre Dame Male/Female Accuracy Dataset (ND-MFAD) [30, 1, 4, 25]. This dataset has several advantages. Images were acquired in a controlled environment over a number of different sessions, so that a gallery of mugshot-quality images can be created. Also, all subjects completed IRB-approved informed consent for their images to be used in research. For our experiments, the images were cropped and aligned using img2pose [2]. The ND-MFAD dataset contains images for 575 subjects who identified as Caucasian female, with 6,669 images total, and for 687 subjects who identified as Caucasian male, totaling 8,419 images.
Figure 5: Cosine similarity distributions comparing images with smiling expressions to those with real and synthetic sunglasses. The observed shift towards higher similarity scores with synthetic sunglasses, compared to real ones, shows that the FaceLab app’s sunglasses addition can preserve identity.
Figure 6: Example of probe images for six conditions considered in our work. (a) original image, (b) image with sunglasses, (c) blur ($\sigma=4.6$), that degrades accuracy about equal to sunglasses, (d) low resolution (37x37), that degrades accuracy about equal to sunglasses, (e) sunglasses + blur (f) sunglasses + low resolution.
In this study, we used a commercial app, “FaceLab [^2] https://lyrebirdstudio.net/,” to add sunglasses to each original image. Figure 2 illustrates the pipeline of adding sunglasses to the face images. Our experimental procedure involved processing 15,088 images through FaceLab app using an Android emulator. Occasional technical issues, including emulator crashes and processing failures, resulted in an initial 2% to 3% failure rate. Any failed or incorrect image processing instances were manually processed to have a “sunglasses-added” version for all the original images. Examples of synthetic sunglasses images by FaceLab app are shown in Figure 3.
While the AR dataset is not appropriate for one-to-many identification experiments, it does contain a subset of images featuring subjects wearing real sunglasses. This characteristic allows us to evaluate the identity preservation capability of the FaceLab app when synthetically adding sunglasses to images. To conduct this assessment, we applied synthetic sunglasses to the normal expression images from the AR database using FaceLab. Examples of image comparisons are shown in Figure 4. We then analyzed the cosine similarity distributions, shown in Figure 5, between matching (a) image of a person smiling to image of the person wearing real sunglasses, and (b) same image of a person smiling to image of the person with synthetic sunglasses added. The results show that the distribution of similarity scores with synthetic sunglasses is shifted to higher similarity compared to the distribution with real sunglasses. It is possible that the strong specular highlights in the images with real sunglasses provide an “extra” lowering of the similarity score. In any case, these results provide additional confidence that the operation of adding sunglasses to any image using the FaceLab app is an identity-preserving operation.
The resulting dataset, named “ND-sunglasses”, allows us to conduct experiments where the addition of sunglasses is the sole different between two probe images of a person. To the best of our knowledge, this dataset represents the largest available collection of paired images with and without sunglasses, specifically curated for experiments on one-to-many facial identification.
IV Experiments
The baseline experiment is as follows. For each person in the dataset, their most recent image is taken as their probe image and their older images are their enrolled images for the gallery. For each probe image, we compute its rank-one mated similarity score and rank-one non-mated similarity score. The mated similarity score is the highest similarity between the person’s probe image and one of the same person’s enrolled images. The non-mated similarity score is the highest similarity between the person’s probe image and an enrolled image of a different person. If the non-mated similarity score is higher than the mated similarity score, then this probe image results in a rank-one false positive identification. To consider results across a set of persons, we look at the distributions of mated and non-mated similarity scores. The more separated these distributions are, the generally greater the accuracy of one-to-many identification. To directly focus on the frequency of false positive identifications, we also consider the distribution of (mated – non-mated) scores across a set of probes. Each negative value in this distribution is from a probe that generates a false positive identification. The fraction of such probes gives the false positive identification rate.
Figure 7: Scores of differences (mated - non-mated) distributions for original probe images and different conditions of probes (top to bottom): 1) sunglasses, 2) blur of $\sigma=4.6$, 3) 37x37 resolution, 4) sunglasses + blur, 5) sunglasses + 37x37 resolution. The face matcher is AdaFace.
ArcFace AdaFace
CF CM CF CM
Effects added to probe mated non-mated mated non-mated mated non-mated mated non-mated
Sunglasses 2.6389 0.3597 2.8259 0.4723 2.6694 0.3003 2.8283 0.4357
Blur ($\sigma=4.6$) 2.7627 0.1653 2.7681 0.2128 2.8102 0.2099 2.7823 0.1506
Low resolution (37x37) 2.6253 0.2104 2.7504 0.2569 2.7011 0.0243 3.2738 0.0686
Sunglasses + Blur
5.2697 0.8537 5.1398 0.5963 5.7830 0.7960 5.1496 0.6423
Sunglasses + Low resolution
4.4400 0.4554 4.5986 0.5141 5.2697 0.8537 4.5865 0.5505
TABLE I: The d-prime values are from rank-one similarity distributions. Similar d-prime values are highlighted in Purple. The data demonstrates that adding sunglasses, introducing blur of $\sigma=4.6$, and reducing resolution to 37x37 in probe images have comparable effects on accuracy. This is notable because, although adding sunglasses to an image might appear to maintain high quality to human observers, it impacts the system’s performance similarly to more obvious image degradation.
The baseline experiment is done with the original mugshot-quality images for probes and enrolled gallery. We then have a series of comparison experiments. Each comparison experiment involves using a modified version of the probe images, matched against the same original mugshot-quality enrolled gallery. For example, the core comparison is made by adding sunglasses to each of the probe images, and observing how the accuracy of one-to-many identification changes compared to the baseline experiment. Making a comparison to the results of the baseline experiment can be done in several ways. One, we can observe the difference in the distributions of the mated scores, non-mated scores, or (mated – non-mated) scores. To give a quantitative estimate to the difference in distributions, we compute the d-prime or the Wasserstein distance [14] between distributions. For example, we can compute the d-prime (or Wasserstein) between the baseline (mated – non-mated) distribution and the (mated – non-mated) distribution for the probe images with sunglasses. A larger d-prime (or Wasserstein) means that sunglasses had a larger effect in degrading accuracy.
The d-prime is commonly used as a quantitative measure of the difference in distributions in biometrics [33, 32, 24]. However, d-prime assumes that the distributions are approximately Gaussian, and the Wasserstein distance (also called “Earth Mover’s Distance) does not. The Wasserstein distance is not currently commonly used in comparing distributions in biometrics, but could be a better choice especially when the distributions are distinctly not Gaussian.
To consider possible accuracy differences across gender, we compute results separately for male and female. For fair comparison across gender, because one-to-many accuracy is affected by the number of enrolled persons and images [24], and there are more males than females in the dataset, we randomly selected 575 Caucasian males to make the number of identities the same for male and female.
We train two recognition models with the ResNet100 backbone and the Glint360K training set [3]. One is trained with ArcFace loss [6] and one with AdaFace loss [17]. The pattern of accuracy loss due to sunglasses in the probe image is similar between the two matchers. Due to space limits, some ArcFace results are in Supplementary Material.
IV-A Accuracy Degradation by Sunglasses in Probe
This section compares one-to-many identification accuracy in two conditions: (a) original high-quality probe without sunglasses matched against high-quality gallery without sunglasses, and (b) original high-quality probe image with sunglasses added matched against same gallery as in (a). Unless otherwise specified, (a) is the baseline.
The mated and non-mated distributions for the two demographic groups are shown in the supplementary material. The d-prime values derived from these distributions are presented in Table I. In the Adaface results, the d-prime values of mated distributions, 2.6694 for females and 2.8283 for males, suggest a more pronounced effect on males. The d-prime values of non-mated distributions, 0.3003 for females and 0.4357 for males, show that the primary effect is in the mated distributions for probes with sunglasses, both female and male. Due to the quality of the original images and the size of the dataset, the mated and non-mated distributions for probes with sunglasses are still separable, but they have moved closer together. This shows that the presence of sunglasses in a high-quality probe significantly lowers accuracy. And, in a real surveillance scenario, images captured from surveillance video often have lower resolution and/or an increased level of blur. Therefore, we explore the effect of combining blur and lower resolution with sunglasses.
IV-B Effect of Sunglasses + Blur / Resolution
Since [24] comprehensively analyzed the effect of the probe having different degrees of blur and reduced resolution in one-to-many identification experiments, we selected a degree of blur and lower resolution that causes a similar impact on accuracy as sunglasses. Table I shows that when adding $\sigma=4.6$ Gaussian blur to the original probe image or resizing the probe image to 37x37, we obtain a similar d-prime value for the mated distribution as the probe having sunglasses. These were then selected as representative blur and resolution values to investigate combined effects of degraded quality. The image quality in actual surveillance scenarios can, of course, vary widely.
Effects added to probe Demographic Distance
Sunglasses CF 0.191
CM 0.190
Blur ($\sigma=4.6$) CF 0.194
CM 0.190
37x37 resolution CF 0.188
CM 0.187
Sunglasses + blur CF 0.370
CM 0.360
Sunglasses + Low resolution CF 0.336
CM 0.333
TABLE II: Wasserstein distance for shift from baseline caused by effects in probe image. Distributions in Figure 7. Face matcher is AdaFace.
Effects added to probe AdaFace ArcFace
CF CM CF CM
None (Original) 0 0 0 0
Sunglasses 0.522 0 0 0
Blur ($\sigma=4.6$) 0.174 0 0 0
37x37 resolution 0 0 0 0
Sunglasses + blur 4.522 1.565 3.652 0.870
Sunglasses + Low resolution
3.304 0.870 3.652 0.696
TABLE III: FPIR for one-to-many identification experiments under various effects in the probe image. As an example, with sunglasses + blur in the probe images, 4.522% of Caucasian females probe images had a rank-one false positive identification with AdaFace.
Figure 6 shows one example of probe images for six conditions: 1) Original high-quality image, 2) Sunglasses added to original, 3) $\sigma=4.6$ Gaussian blur added to original, 4) Original resized to 37x37, 5) Sunglasses and blur, and 6) Sunglasses and reduced resolution. Note that adding sunglasses to an original image may subjectively appear to result in a high-quality image, even though accuracy is degraded similar to a noticeable degree of blur or low resolution. Table I shows that when combining blur and low resolution with sunglasses, the d-prime values are almost doubled between mated distributions and almost tripled between non-mated distributions, compared with the single factor, for both face matchers and demographic groups. For example, in the AdaFace results for mated females, the d-prime values are 2.6694 when only adding sunglasses to the probes and 5.783 when combining sunglasses and blur in the probes.
IV-C Analysis of Effect on Model Performance
To quantitatively analyze the false positive identifications (FPI) caused by each condition, we report the Wasserstein Distance of the score difference distribution and the FPIR. For a given probe image, a negative score difference indicates an FPI, as the highest similarity value between the probe image and an image from a different person is higher than the top matched image from the same identity. Thus, the distribution of score differences provides a measure of the FPI potential. The distributions of these differences are shown in Figure 7. A general conclusion is that each of the probe quality factors – wearing sunglasses, increased blur, lower resolution – causes the (mated - non-mated) score distribution to shift to lower values, representing increased chances for FPI. The greatest shifts naturally occur when probe images have combined blur and sunglasses, or combined lower-resolution and sunglasses.
The Wasserstein Distance [14] measures the distance between two distributions without assuming that they are Gaussian. It can be used to compare the size of effect of different probe quality factors on the distribution of (mated - non-mated) scores, with larger values indicating larger degradation in accuracy. Table II shows the Wasserstein Distance between the baseline (original vs. gallery) and each of five instances of degraded probe quality. Each of the three individual quality factors (wearing sunglasses, increased blur, decreased resolution) has a similar negative impact on accuracy. This is to be expected, as the particular blur and resolution values were selected for this. The sunglasses + blur degradation is close to the sum of the two individual effects, and sunglasses + lower-resolution degradation is slightly below the sum of the two individual effects. The degree of accuracy degradation is similar for males and females, although females had lower baseline accuracy. The overall message from these results is that accuracy degrades rapidly with a combination of different quality factors.
Conditions CF CM
mated non-mated mated non-mated
One img. per ID 0.5344 0.7142 0.6911 0.8910
All gallery img. 1.3203 1.6576 1.6190 1.8357
TABLE IV: The d-prime values of rank-one similarity distributions for probe images with sunglasses under different gallery conditions: the first row shows sunglasses randomly added to one image per identity in the gallery, the second row shows sunglasses added to all gallery images. The face matcher is AdaFace.
We also calculated the False Positive Identification Rate (FPIR) for each probe quality condition, using both the ArcFace and AdaFace matchers. These results in Table III show that both ArcFace and AdaFace have zero FPIR for the original and 37x37 resolution probes. AdaFace experiences an FPIR of 0.522% for sunglasses-added probes and an FPIR of 0.174% for blurred probes. But the FPIR increases sharply when sunglasses are combined with either blur or reduced resolution. For example, sunglasses + blur results in FPIR of 3.652% for females for ArcFace, and 4.522% for AdaFace. The absolute level of FPIR varies with the size of a dataset, with a larger number of persons enrolled in the gallery causing an increased FPIR [24]. So the overall low FPIR in these results does not project what can be experienced in practice for scenarios with larger galleries. However, the important general point illustrated in these results is that multiple degraded quality factors in the probe image causes a dramatic increase in the FPIR. Also, both matchers consistently show higher FPIR for females than for males. This gender-based difference illustrates another general lesson for one-to-many identification accuracy.
Shift in mated - non-mated
compared to baseline
% of lost accuracy recovered
Experiment Conditions AdaFace ArcFace AdaFace ArcFace
CF 0.191 0.197 - -
sunglasses probe vs. original gallery CM 0.190 0.194 - -
CF 0.171 0.176 10.47% 10.66%
sunglasses probe vs. add sunglasses to one image per ID in the gallery CM 0.164 0.163 13.68% 15.98%
CF 0.136 0.141 28.80% 28.43%
sunglasses probe vs. add sunglasses to all gallery images CM 0.122 0.121 35.79% 37.63%
TABLE V: Wasserstein Distance of the differences in (mated - non-mated) distributions. Adding sunglasses to one image per ID in the gallery recovered 10.47% with AdaFace and 10.66% with ArcFace Blue of the accuracy lost for females and 13.68% with AdaFace and 15.98% with ArcFace Purple for males. Adding sunglasses to all gallery images recovered 28.80% with AdaFace and 28.43% with ArcFace Green of the accuracy lost for females and 35.79% with AdaFace and 37.63% with ArcFace Orange for males.
IV-D Strategy for Recovering Accuracy Loss
Previous sections show that a probe image with sunglasses matched against a gallery of images without sunglasses results in a substantial decrease in accuracy. Primarily, the occlusion of facial features by sunglasses in the probe image obscures vital identifying information. But also, comparing images with and without sunglasses presents a mismatch in facial features, which is potentially a different factor affecting accuracy. One could imagine an approach to remove sunglasses from the probe image, or an approach to add sunglasses to the gallery images. Removing sunglasses from the probe could perhaps require less computation, as it involves the one image, but it requires hallucinating face feature content that is unseen. Adding sunglasses to the gallery images, on the other hand, involves adding a plausible instance of a specific style of occlusion to each image.
In this section, we investigate whether adding sunglasses to the gallery images can mitigate the accuracy degradation that would otherwise occur when the probe has sunglasses.
Table IV The table illustrates the d-prime values of rank-one similarity distributions for probe images with sunglasses under different gallery conditions. When sunglasses are randomly added to a single image per identity in the gallery, with an average of 11.15 images per identity, the results show d-prime values of 0.5344 (mated) and 0.7142 (non-mated) for females, and 0.6911 (mated) and 0.891 (non-mated) for males. The d-prime values become more pronounced when sunglasses are added to all gallery images, as evidenced by the increased values: 1.3203 (mated) and 1.6576 (non-mated) for females, and 1.619 (mated) and 1.8357 (non-mated) for males. Notably, in both scenarios, the improvement in accuracy is more substantial for males than for females. The results demonstrate that under both gallery conditions, accuracy improves; however, adding sunglasses to all gallery images leads to a higher recovery in accuracy. The value of adding sunglasses to the gallery images to ameliorate the accuracy degradation is more obvious in the (mated - non-mated) distributions shown in Figure 8.
Table V presents the Wasserstein Distances of the difference distributions. When we randomly select one image per identity in the gallery to add sunglasses, 1150 images in total, it recovers 10.47% with AdaFace and 10.66% with ArcFace of the lost accuracy for females and 13.68% with AdaFace and 15.98% with ArcFace for males when the person in the probe image is wearing sunglasses. Furthermore, when sunglasses are added to all gallery images, we observe a more substantial recovery of 28.8% with AdaFace and 28.43% with ArcFace of the lost accuracy for females and 35.79% with AdaFace and 37.63% with ArcFace for males. It’s worth noting that both the d-prime and Wasserstein Distance metrics showed similar general patterns in our results. However, visual inspection suggests the distributions are not strictly Gaussian. Given this, the Wasserstein Distance may be a more appropriate metric for this type of study, even though in this case the overall pattern of results was not substantially different between the two measures.
It is important to note that 100% recovery of lost accuracy is not feasible due to the occlusion of facial features by sunglasses and because the type of sunglasses in the gallery is not matched to those in the probe, so even for the same identity, probe and gallery images may contain different sunglasses. Nevertheless, our results show the potential of this approach to recover part of the lost accuracy without re-training the models, thereby enhancing the robustness of one-to-many facial identification systems in scenarios involving occlusions. In a real-world scenario, it may be more advantageous to consistently add sunglasses similar to those encountered in probe images to the gallery images.
Figure 8: Similarity difference (mated - non-mated) distributions. Comparing with the baseline, adding sunglasses to probe images causes a large accuracy loss. To counter this effect, we add sunglasses to one (all) image(s) per identity in the gallery to achieve good accuracy recovery, as shown in Top (Bottom) row. The face matcher is AdaFace.
IV-E Effect of Sunglasses Training Samples in Accuracy
Accuracy lost by adding sunglasses to face can possibly be explained by the under-representation of wearing-sunglasses samples in the training set. Thus, we start the analysis on a widely recognized training set, WebFace4M [40], which encompasses 4M images from 200K identities.
Considering the absence of dark sunglasses labels of the WebFace4M training set and off-the-shelf sunglasses classifiers, we developed a specialized dark sunglasses detector to obtain the number of wearing-sunglasses samples in WebFace4M. Specifically, we first use a facial attribute classifier [36] to get the wearing-eyeglasses images in WebFace4M. Part of these images are then manually separated to three classes: no-eyeglasses, eyeglasses, and sunglasses, for the next training step. This process results in 8,500 dark sunglasses images, 8,200 eyeglasses images, and 8,200 no-eyeglasses images. This dataset is subsequently partitioned into training, validation, and test sets, with allocations of 70%, 15%, and 15%, respectively. We then fine-tuned a ResNet101 [12], which is pre-trained on ImageNet, using this dataset. This dark sunglasses detector achieves 93.3% accuracy on the test set. After running this detector on WebFace4M, there are 115,818 (3.8%) images are predicted as having dark sunglasses, showcasing the wearing-sunglasses samples are under-represented in the training set. Hence, increasing the representation of wearing-sunglasses images in the training set is a possible way to improve the accuracy.
Sunglasses Types sunglasses images (%)
0% 10% 20% 23%
real sunglasses 95.31% 95.18% 95.05% 95.22%
synthetic sunglasses 95.31% 95.24% 95.11% 95.17%
TABLE VI: A general face recognition test accuracy (%) on real and synthetic sunglasses datasets across models trained with varying proportions (0%, 10%, 20%, and 23%) of sunglasses-augmented images.
However, adding sunglasses images to the dataset is non-trivial, so we investigate how the fraction of sunglasses images impacts the accuracy. First, we downscale the dataset to 500K by randomly selecting the no-sunglasses images from 38,802 identities that are detected having sunglasses images. In this scale, the detected wearing-sunglasses images occupy roughly 23% of the dataset. We then train the models with varying proportions of wearing-sunglasses images by replacing the no-sunglasses images with wearing-sunglasses images. These models are tested on ND-Sunglasses to obtain the FPIR trend w.r.t. different proportions. The left plot in Figure 9 indicates a clear decreasing FPIR when increasing the fraction of sunglasses images. From 0% to 23%, the FPIR has a 42.87% decrease for CF and all CM probe images are correctly identified.
To efficiently add new wearing-sunglasses images to the dataset, we choose to use a face synthesizing model, Vec2Face [34]. It can effectively generate the images of the same identity with sunglasses, as shown in Figure 10. We then conduct the same experiments to evaluate if the synthesized wearing-sunglasses images can achieve the similar effect to the real ones. Note that, we uniformly generate images for all 38,082 identities. The right plot in Figure 9 indicates that using synthesized images can achieve better performance than real ones. The FPIR has a 57.17% decrease for CF and all CM probe images are correctly identified. This showcases the possible accuracy improvement in the full-size WebFace4M by adding synthesized wearing-sunglasses images in the future work. It is worth noting that all the trained models have close average accuracy on five standard test sets [15, 38, 39, 23, 27], as shown in Table VI.
Figure 9: This illustrates FPIR changes as the proportion of sunglasses-augmented images in the training set from 0% to 23%, for both real (left) and synthetic (right) sunglasses. Performance is evaluated in one-to-many identification experiments where probe images are wearing sunglasses.
Figure 10: Examples of original images (top row) and generated images with sunglasses (bottom row) using the Vec2Face generative model.
V Conclusions and Future Work
This study systematically investigates the impact of sunglasses in the probe images on one-to-many identification accuracy, addressing a gap in existing literature. We assembled a dataset, ND-Sunglasses, to support the investigation. Our findings reveal that sunglasses in probe images substantially reduces the accuracy in one-to-many identification, comparable to applying Gaussian blur with $\sigma=4.6$ or reducing image resolution to 37x37 pixels, despite these latter factors causing more perceptible visual changes.
Our research also demonstrates that increasing the proportion of sunglasses-wearing images in the gallery can partially mitigate the accuracy loss caused by sunglasses in probe images. Specifically, adding sunglasses to all gallery images recovered 35.79% of lost accuracy for females and 37.63% for males with ArcFace.
We show that the accuracy degradation caused by wearing sunglasses is due to the under-representation of wearing-sunglasses images. Moreover, we show that using a face synthesizing model to uniformly add wearing-sunglasses images to the dataset can reduce the error rate. Future work is to deploy the same strategy on a full-size training set.
VI Ethical Impact Statement
Documented instances of wrongful arrest have brought critical attention to one-to-many facial identification. This research improves understanding how quality factors in a probe image degrade accuracy, and should lead to more realistic appraisal of the accuracy of one-to-many identification, hopefully contributing to reducing instances of wrongful arrest.